# Adaptive Group Policy Optimization: Towards Stable Training and Token-Efficient Reasoning

### v2

@@ Abstract
Since DeepSeek-R1 popularized, Group Relative Policy Optimization (GRPO) has become the core part of training Reasoning LLMs. However, we find some deficiency that influences RL stability and inference efficiency, like zero-variance in advantage estimation. Thus, we propose Adaptive Group Policy Optimization (AGPO) which contains a simple but effective modification: a revised objective function to mitigate training fluctuation and zero advantage. The experiments demonstrate our method achieves more stable training and superior performance with significantly fewer tokens in reasoning steps.
@@ 1 Introduction
For alignment challenges, reinforcement learning from human feedback (RLHF)Bai et al. (2022); Kaufmann et al. (2023) is introduced as a prominent post-training strategy, adopted by notable LLMs including GPT-4, Claude, Gemini, and DeepSeek. They have explored various optimization techniques such as Proximal Policy Optimization (PPO) Schulman et al. (2017) and Direct Preference Optimization (DPO) Rafailov et al. (2023). Recently, to significantly reduce computational and memory overhead associated with PPO, DeepSeek eliminated the value model and proposed Group Relative Policy Optimization (GRPO)Guo et al. (2025), which achieved high computational efficiency and excellent reasoning performance, surpassing other open-source models ranging from 7B to 70B.
@@
(1) Confusing Training Signal: Negative losses happen in RL training, but in this scenario it is not always beneficial. Higher group accuracy may have lower advantage thus influencing loss estimation. Besides, when all rewards within a group are identical, the normalized advantage approaches 0, causing the loss signal to vanish, which potentially stalling training.
(2) Inefficient CoT Length: Since GRPO lacks mechanisms to discourage excessively long chain-of-thought (CoT), models tend to produce overly verbose explanations. A refined approach that rewards concise and effective reasoning is essential to improve token efficiency.
To address these issues, we propose an enhanced GRPO training algorithm, Adaptive Group Policy Optimization (AGPO). Our main contributions are summarized as follows:
  • Training Efficiency: By identifying the limitations of GRPO’s advantage, we introduce an adaptive loss function that addresses negative loss and zero advantage scenarios, ensuring continuous and effective learning.
@@ 2.3 Group Relative Policy Optimization (GRPO)
$J_{\text{GRPO}}(\theta)&=\mathbb{E}_{(q)\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\Bigg{(}\min\Bigg{(}\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}{A}_{i},\\ &\quad\quad\quad\operatorname{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},1-\epsilon,1+\epsilon\right){A}_{i}\Bigg{)}-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\Bigg{)}\Bigg{]}\end{aligned}$
@@ 3 Adaptive Group Policy Optimization
During RL training with the GRPO algorithm, we find a useful modification to improve performance and efficiency.
3.1 Adaptive Loss
It is easy to find that if the rewards of the group are all equal, like all correct or all wrong, then the advantages become all zero. Therefore, the corresponding sample in the batch has no effect in the training. As model training progresses, this situation increases. We believe that it is helpful to mask out these positions from the loss calculation.
Another issue is entropy collapse occurred in GRPO training. We observed the same phenomenon as in other works, like DAPO (Yu et al., 2025). The entropy drops quickly leading the policy model to give up exploring. However, we consider it as the problem of objective design rather than clip ratio. Current negative loss value tends to be overconfident since a group with higher accuracy may have higher deviation of rewards which may cause lower advantage compared to other groups in batch. We propose an adaptive loss to tackle these issues:
@@ 3.1 Adaptive Loss
Usually in RL training, the model is updated through a mini batch size of training data. Therefore, the loss mask in Equation 6 takes place in the mean operation of losses of questions $q$ in the batch. Negative loss values are also clipped for the mean loss of the batch in order to control fluctuation and maintain a suitable training level. We believe the clip avoids the model getting stuck in a local optimum.
By replacing the original objective, our method focuses on useful information in the batch and normalizes the loss for stable training, which would bring performance improvement and token efficiency.
4 Experiment
@@ 4.1 Implementation Details
We use Qwen2.5-7B and Qwen2.5-14B as the base models. All experiments are conducted on our curated dataset, which is constructed by mixing data from MATH train set (Hendrycks et al., 2021) and DAPO train set (Yu et al., 2025). For MATH train set, only data where difficulty levels are greater than or equal to 3 are selected. Similarly, for DAPO train set, data are retained only if solution rates achieved by Qwen3-32B model (Yang et al., 2025) are fall between 0.5 and 0.8 inclusively. These filtering techniques are to ensure the difficulty distribution across the obtained dataset is balanced.
VeRL (Sheng et al., 2024) is utilized to perform RL training with a train batch size of 32, a PPO mini batch size of 8 and a learning rate of $61e-$ 6. The number of group rollout is 8. Temperature for generation is set to 1. As for reward settings, the $r_{correct}$ and $r_{wrong}$ for accuracy reward are set to 0 and 1 respectively. Trained checkpoints that achieve best performance on MATH-500 (Lightman et al., 2023) with the metric of Pass@1 are selected for further evaluation with respect to token efficiency.
It is worth noting that KL divergence penalty is not applied for all experiments. This is based on the observation that model distribution can vary significantly compared to reference model during long CoT training. Therefore, removing KL divergence has been adopted as common practice in the domain.
@@ 4.2 Main Results
Table 1 shows the performance of different models on the benchmark. Both GRPO and AGPO acquire huge performance gains compared with the base models. As for Qwen2.5-7B experiments, we observe a clear improvement on MATH-500 from 73.2 to 74.6 at the best checkpoint. Qwen2.5-14B experiments also induce similar conclusion that the adaptive loss further refines the model by 1.8 percentage. It is obvious that both loss clip and loss mask are important for our method. If we train without the mask, the performance drops even lower than that of GRPO for all model sizes and it is same for loss clip.
Table 2 illustrates the token efficiency of different models on the benchmark. Qwen2.5-7B-AGPO only takes 533 tokens on average for solving MATH-500 problems while Qwen2.5-7B-GRPO consumes 699 tokens that is $31\%$ higher. In terms of 14B models, our AGPO also ranks first in token efficiency which uses 521 tokens averagely.
@@ 4.3 Training Dynamics
We examine several variations of training metrics after applying AGPO.
As shown in Figure 1 and Figure 2, a significant enhancement in entropy is observed for both Qwen2.5-7B and Qwen2.5-14B actor models after the application of AGPO. This observation can be attributed to the lower loss clip operation in AGPO, where only positive training loss from below-average actions is maintained while negative training loss is overconfident and controlled in the gradient update. Consequently, the probability distribution is drifted upward asymmetrically by adequate gradient without being drifted downward meanwhile, which manifests higher measured entropy during training. The higher entropy eventually facilitates the generation of more diversified samples within the batch, which is essential for large-scale RL training.
@@
We also find substantial reductions in response length while the training steps increase, for both models with AGPO as Figure 3 and Figure 4 show. Meanwhile, comparable accuracy performance is maintained on training set as shown in Figure 5 and Figure 6. This phenomenon can be attributed to masking loss operation in AGPO, since the effect of loss mask on response length is clearly shown in Figure 3 and Figure 4. In this operation, the losses of all correct or wrong groups are masked, enhancing accurate estimation of average loss of the batch. The correction consequently amplifies gradient update towards correct direction. Therefore, in combination with the empirical observation, we believe the loss mask implicitly serves as length-based reward to constrain response length.
@@ 5 Conclusion
In this work, we propose a novel method, AGPO, to train a more powerful reasoning model. Our adaptive loss, including loss clip and mask, demonstrates noticeable improvement on both model performance and inference efficiency. Also, our method helps to avoid entropy collapse while training. We want to do more designs about the adaptive loss as future directions. For example, the current loss can be normalized to non-negative values by exponential equations.

### v3

@@ Abstract
Since DeepSeek-R1 popularized, Group Relative Policy Optimization (GRPO) has become the core part of training Reasoning LLMs. However, we find some deficiency that influences RL stability and inference efficiency, like zero-variance in advantage estimation. Thus, we propose Adaptive Group Policy Optimization (AGPO) which uses a simple but effective method, an adaptive loss function, to mitigate training fluctuation and token inefficiency. The experiments demonstrate our method achieves more stable training and superior performance with significantly fewer tokens in reasoning steps.
@@ 1 Introduction
For alignment challenges, reinforcement learning from human feedback (RLHF)Kaufmann et al. (2023) is introduced as a prominent post-training strategy, adopted by notable LLMs including GPT-4, Claude, Gemini, and DeepSeek. They have explored various optimization techniques such as Proximal Policy Optimization (PPO) Schulman et al. (2017) and Direct Preference Optimization (DPO) Rafailov et al. (2023). Recently, to significantly reduce computational and memory overhead associated with PPO, DeepSeek eliminated the value model and proposed Group Relative Policy Optimization (GRPO)Guo et al. (2025), which achieved high computational efficiency and excellent reasoning performance, surpassing other open-source models ranging from 7B to 70B.
@@
Confusing Training Signal: When all rewards within a group are identical, the normalized advantage approaches 0, causing the loss signal to vanish, which potentially stalling training. Besides, negative losses happen in RL training, but in this scenario it is not always beneficial.
Inefficient CoT Length: Since GRPO lacks mechanisms to discourage excessively long chain-of-thought (CoT), models tend to produce overly verbose explanations. A refined approach that rewards concise and effective reasoning is essential to improve token efficiency.
To address these issues, we propose an enhanced training algorithm, Adaptive Group Policy Optimization (AGPO). Our main contributions are summarized as follows:
  • Training Stability: By identifying the limitations of GRPO’s advantage, we introduce an adaptive loss function that addresses the scenarios of zero advantage and negative losses, ensuring continuous and effective learning.
@@ 2.3 Group Relative Policy Optimization (GRPO)
$\begin{aligned} J_{\text{GRPO}}(\theta)&=\mathbb{E}_{(q)\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\Bigg(\min\Bigg(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}{A}_{i},\\ &\quad\quad\quad\operatorname{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},1-\epsilon,1+\epsilon\right){A}_{i}\Bigg)-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\Bigg)\Bigg]\end{aligned}$
@@ 3 Adaptive Group Policy Optimization
Performance fluctuation and entropy collapse always happen in the later stage of GRPO training, and we want to tackle it from the perspective of loss function. The loss function of GRPO has two issues:
  • Zero advantage and thus zero loss: It is caused by groups with all-correct or all-wrong responses. This is not effectively being ignored during optimization by GRPO. According to Equation 5, the mean loss of a batch is used for model update, and zero losses in the batch lead to misestimation of the scale.
  • Positive advantage and thus negative loss: We observed entropy collapse same as other works like DAPO (Yu et al., 2025). The entropy drops quickly leading the policy model to give up exploring. However, we consider it as the problem of negative loss rather than clip ratio. It makes the model overconfident and tends to ignore incorrect samples in the batch.
Therefore, we propose an adaptive loss to tackle these issues as shown in Equation 6. By replacing the original objective, our method focuses on useful information in the batch and normalizes the loss for stable training, which brings performance improvement and token efficiency.
@@
3.1 Loss Mask
It is easy to find that if the response rewards are all equal in a group, like all correct or all wrong, then the advantages become all zero. The phenomenon of zero advantages naturally becomes more often during training as the model becomes better. Usually in RL training, the model is updated through a mini batch of training data. The corresponding groups in the batch have no effect in the training. Loss mask does not aim to reduce the percentage of such groups, but to avoid including them in the loss computation. Therefore, the loss mask in the first part of Equation 6 means excluding invalid samples in the mean operation of sample losses.
3.2 Loss Clip
Clipping negative batch loss to zero is a simple but effective way to penalize the overconfidence, which achieves solid performance improvements. Another work (Zhu et al., 2025) has similar conclusion that training with only negative samples can be highly effective.
4 Experiments
@@ 4.1 Implementation Details
We use Qwen2.5-7B and Qwen2.5-14B as the base models. All experiments are conducted on our curated dataset, which is constructed by mixing data from MATH train set (Hendrycks et al., 2021) and DAPO train set (Yu et al., 2025). We remove overly easy and overly hard data points from these sets. Specifically, for MATH train set, only data where difficulty levels are greater than 2 are selected. Similarly, for DAPO train set, data are retained only if solution rates achieved by Qwen3-32B model (Yang et al., 2025) fell between 0.5 and 0.8 inclusively.
VeRL (Sheng et al., 2024) is utilized to perform RL training with a train batch size of 32, a PPO mini batch size of 8 and a learning rate of 1e61e-6. The number of group rollout is 8. Temperature for generation is set to 1. As for reward settings, the accuracy rewards $r_{correct}$ and $r_{wrong}$ are set to 0 and 1 respectively. Checkpoints that achieve best performance on MATH-500 (Lightman et al., 2023) with the metric of Pass@1 are selected for further evaluation with respect to token efficiency.
It is worth noting that KL divergence penalty is not applied for all experiments. This is based on the observation that model distribution can vary significantly compared to reference model during long CoT training. Therefore, removing KL divergence has been adopted as common practice in the domain (Schulman et al., 2017; Yu et al., 2025).
@@ 4.2 Main Results
Table 1 shows the performance of different models on the benchmark. Both GRPO and AGPO acquire huge performance gains compared with the base models. As for Qwen2.5-7B experiments, we observe a clear improvement on MATH-500 from 73.2 to 74.6 at the best checkpoint. Qwen2.5-14B experiments also give same conclusion that AGPO further refines the model by 1.8 percentage. It is obvious that both loss mask and loss clip are important for our method. If we train without the mask, the performance drops even lower than that of GRPO for all model sizes and it is same for loss clip.
Table 2 illustrates the token efficiency of different models on the benchmark. Qwen2.5-7B-AGPO only takes 533 tokens on average for solving MATH-500 problems while Qwen2.5-7B-GRPO consumes 699 tokens that is $31\%$ higher. In terms of 14B models, our AGPO also achieves best in token efficiency which uses 521 tokens averagely.
@@ 4.3 Training Dynamics
We examine several training metrics of AGPO for further analysis.
As shown in Figure 1 and Figure 2, a significant enhancement to entropy is observed for both Qwen2.5-7B and Qwen2.5-14B actor models compared with GRPO. This observation can be attributed to the loss clip operation in AGPO, where only positive training loss from below-average actions is maintained while negative training loss is controlled in the gradient update. Consequently, the probability distribution is drifted upward asymmetrically by adequate gradient without being drifted downward meanwhile, which manifests higher measured entropy during training. The higher entropy eventually facilitates the generation of more diversified samples within the batch, which is essential for large-scale RL training.
@@
Figure 3 and Figure 4 indicate that substantial reductions in response length are achieved for both AGPO models. Meanwhile, comparable accuracy performance is maintained on training set as shown in Figure 5 and Figure 6. We believe the loss mask implicitly serves as length-based reward to constrain response length, since the response length of AGPO is significantly shorter than that of AGPO (w/o loss mask). We also observe from the figures that the response length for AGPO is similar to GRPO at the early training stage, but eventually becomes shorter at middle and final stages compared with the GRPO baseline. This is because the initial model parameters are similar for both AGPO and GRPO. However, by masking all correct and all wrong groups where advantages are zero, the loss distribution within the batch is re-normalized as training progresses. This amplifies the gradient updates for AGPO so it tends to encourage the token distribution of correct response with higher confidence. Hence, AGPO requires fewer tokens or thinking words to arrive at the correct response, whereas GRPO needs more training iterations and response tokens to approach similar results.
4.4 Generalization
We conduct a set of additional experiments on code reasoning for assessing generalization of our method. Deepseek-R1-Distill-Qwen-7B is chosen as the base model for post RL training. We also curate 8.4K training data from CodeForce (Penedo et al., 2025) and CodeContest (Li et al., 2022). We use LiveCodeBench (2024.10-2025.1) (Jain et al., 2024) as the benchmark for evaluation. The results can be seen in Table 3 that our method significantly outperforms GRPO by 3 points. We also observe the performance of GRPO drops dramatically from 40.1 to 23.5 after 40 steps, while the performance of AGPO keeps increasing.
Model LiveCodeBench (Pass@1)
Deepseek-R1-Distill-Qwen-7B 34.3
Deepseek-R1-Distill-Qwen-7B-GRPO 40.1
Deepseek-R1-Distill-Qwen-7B-AGPO 43.1
Table 3: Performance of different RL techniques on LiveCodeBench.
@@ 5 Conclusion
In this work, we propose a novel method, AGPO, to train a more powerful reasoning model. Our adaptive loss, including loss mask and clip, demonstrates noticeable improvement on both model performance and token efficiency. Our method also helps to avoid entropy collapse while training. As for future directions, we want to do more designs about the adaptive loss. For example, the current loss can be normalized to non-negative values by exponential equations.